OpenStudy (anonymous):

in general,when a number written in scientific notation includes aa positive exponent,the magnitude of the number is what?

OpenStudy (anonymous):

the value of the exponent equals the number of digits after the first digit e.g. 10^2 = 100 10^4 = 10.000 2.1 * 10^6 = 2,100,000